Encyclopedia Astronautica
2008.03.23 - ISS On-Orbit Status 03/23/08


Sunday - J/A Flight Day (FD) 13/14.

Ahead: Week 23 of Increment 16. HAPPY EASTER! Congratulations, Shuttle & ISS crews: Five EVAs in a row, all successful. What a great Easter gift!

>>>>Today at ~7:43am EDT, the ISS, specifically its FGB module, completed 53,500 orbits of the Earth, having covered a distance of 2.25 billion kilometers (1.4 billion st.miles) in 3411 days. The 19,300 kg (42,600 lbs) Zarya ('Dawn') was launched on a Russian/Khrunichev Proton from Baikonur over nine years ago (11/20/1998) as the first element of the multi-national space station.<<<<

Crew sleep/wake cycle today: Sleep 4:00am -12:30pm; wake 12:30pm -3:15am tomorrow.

EVA-5 was completed successfully by Bob Behnken & Mike Foreman in 6h 24m, accomplishing all of its objectives.
During the spacewalk, Behnken (EV1) & Foreman (EV2) -

Installed KAUs (Keep-alive Umbilicals) on the OBSS (Orbiter Boom Sensor System);
Stowed the OBSS on the ISS S0/S1 truss segments, powered it up and installed its cover (OBSS will be returned on 1J);
Installed two MISSE-6 PEC (Materials International Space Station Experiment #6 Passive Experiment Containers) in the LWAPA(Light Weight Adapter Plate Assembly) on the Columbus EPF (External Payload Facility) & photographed LWAPA;
Inspected the Starboard SARJ (Solar Alpha Rotary Joint) covers, i.e., divot (depression) on cover 18 & photograph covers 16, 6, 8, 9 & 11 (no debris or divots discovered here);
Clean up worksite & ingress.
Additionally, two get-ahead tasks were completed:

Installation of the five JLP (JEM Experiment Logistics Module Pressurized Section) & keel pin trunnion thermal covers, and
Stowing the general purpose cutters into the Airlock tool box #2.
Official start time of the spacewalk was 4:34pm EDT, about 49 min ahead of timeline, and it ended at 10:36am. Total EVA duration (PET = Phase Elapsed Time) was 6h 2min. It was the 109th spacewalk for ISS assembly & maintenance and the 81st from the station (59 from Quest, 22 from Pirs, 28 from Shuttle) totaling 498h 49min, the 13th for Expedition 16 (totaling 90h 57min) and the 10th so far this year. After today's EVA, a total of 139 spacewalkers (107 NASA astronauts, 21 Russians, and 11 astronauts representing Japan-1, Canada-4, France-1, Germany-2 and Sweden-3) have logged a total of 687h 11min outside the station on building, outfitting and servicing. It was also the 131st spacewalk involving U.S. astronauts.

After crew wakeup at ~12:30pm EDT yesterday, ending the 8.5-hr sleep period, the Airlock Crewlock (A/L CL) hatch was cracked at ~2:00pm for a hygiene break/with mask prebreathe for Behnken & Foreman after spending the night on 10.2 psi campout. Around 2:35pm, the hatch was closed again by Peggy Whitson, the IV (Intravehicular Crewmember) in charge, for EVA preparations in 10.2 psi, followed by EMU purge & prebreathe. Afterwards, the IV supported CL depressurization and EV1/EV2 egress (EMUs switched to batteries at 4:34pm).

As part of pre-EVA activities, FE-1 Malenchenko powered down the ham radio equipment in SM (Service Module) and FGB at ~1:53pm EDT to prevent RF interference with the EMUs during the spacewalk.

FE-1 Yuri Malenchenko began his 'day' by attending to the current experiment session with the Russian/German TEKh-20 Plasma Crystal-3+ (Plazmennyi-Kristall/PK-3+) payload, activating the turbopump in the Service Module (SM)'s Transfer Compartment (PkhO) for keeping the vacuum chamber (ZB) in the SM Work Compartment (RO) evacuated. The turbopump was then deactivated again this morning at ~3:55am EDT before sleeptime. (Main objective of PK-3 is to study dust plasma wave propagation and dispersion ratio at a specified power of HF discharge, pressure, and a varied number of particles.)

CDR Peggy Whitson conducted a thorough CO2 and air exchange survey, measuring the airflow between the Lab and Node-2 IMV (Intermodular Ventilation) using the Velocicalc instrument and reporting the measurements to MCC-Houston for analysis by ground specialists to aid in the CO2 level troubleshooting. (Measurements included CO2 readings in the Lab, Node-2 aft and forward, Shuttle middeck, Shuttle flight deck & Shuttle A/L using the CDMK (Carbon Dioxide Monitor Kit). CO2 data indicated a 1 mmHg partial pressure difference between the US Lab and Node-2 (3.13 & 4.2 mmHg, resp.). The Node-2 atmosphere is a mixture from Node-2, Columbus, JLP and Orbiter, while US Lab air is a mixture from US Lab, Node-1, and A/L. Ground teams have now developed an additional path for routing air, and the crew was given the Go this morning by the IMMT (ISS Mission Management Team) for installing a 23-ft IMV duct from Node-2 to the US Lab to facilitate CO2 removal by the CDRA.)

Later tonight, Whitson will install the 23 feet of flexible ventilation ducts from the Node-2 Forward Stbd End Cone IMV Return (bypassing the noise muffler) to the Node-2 Aft Hatch to aid in CO2 removal. (This configuration will bypass the stbd IMV return of the Node 2/Lab vestibule by returning air directly through the hatch.)

FE-1 Malenchenko and FE-2 Leo Eyharts had several hours reserved for transferring more discarded cargo to the Progress 28P and stowing it according to detailed plan in the logistics spacecraft-turned-trash can, to be jettisoned on 4/7.

The FE-1 performed scheduled IFM (in-flight maintenance) on the SM's condensate water processor (SRV-K2M) by removing and replacing its water-conditioning unit purification columns (BK BKV). The old unit was disposed of in Progress 28P. (Last time done: 9/12/07) (The SRV-K2M converts collected condensate into drinking water and dispenses the reclaimed water (warm and hot) for drinking and preparation of food and beverages.)

Before breakfast and exercise, FE-2 Eyharts performed his second PHS (Periodic Health Status) w/Blood Labs examination. Garrett Reisman assisted in drawing blood and using the U.S. PCBA(Portable Clinical Blood Analyzer). The second part of PHS, Subjective Clinical Evaluation, was performed later in the day. (The PHS exam, with PCBA analysis and clinical evaluation, is guided by special software (IFEP, In-Flight Examination Program) on the MEC (Medical Equipment Computer). While PCBA analyzes total blood composition, the blood's hematocrit is particularly measured by the Russian MO-10 protocol.)

Eyharts also conducted his second session with the biomed experiment INTEGRATED IMMUNE (Validating Procedures for Monitoring Crew member Immune Function), collecting wet saliva samples first thing in post-sleep. (IMMUNE protocol requires the collection to occur first thing post-sleep, before eating, drinking and brushing teeth, and all samples are stored at ambient temperature. Along with NUTRITION (Nutritional Status Assessment), INTEGRATED IMMUNE samples & analyzes participant's blood, urine, and saliva before, during and after flight for changes related to functions like bone metabolism, oxidative damage and immune function to develop and validate an immune monitoring strategy consistent with operational flight requirements and constraints. The strategy uses both long and short duration crewmembers as study subjects. The saliva is collected in two forms, dry and liquid. The liquid saliva collections require that the crewmembers soak a piece of cotton inside their mouth and place it in a salivette bag; there are four of the liquid collections during docked operations.)

FE-2-16 Reisman had his third IMMUNE saliva collection this afternoon after wake-up (~12:30pm), this time dry samples. (The dry samples are collected at intervals during the collection day using a specialized book that contains filter paper, all stored at ambient temperature.)

The FE-1 performed the periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various RS (Russian Segment) hatchways, including the FGB-to-Soyuz tunnel, and the FGB-to-Node passageway. (This is especially important when the ventilation/circulation system has to cope with a larger crew on board, currently ten persons, and one of the two Russian SKV air conditioners off (SKV-1).),

Yuri also completed the routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including ASU toilet facilities systems/replaceables plus the weekly collection of the toilet flush counter (SPK-U), water supply (SVO) readings and POTOK parameters for calldown to TsUP/Moscow. (Regular daily SOZh maintenance consists of replacement of the KTO & KBO solid waste containers, replacement of an EDV-SV waste water and EDV-U urine container, replacement of the KOV EDV for the Elektron-intended water, and processing U.S. condensate water as it becomes available in a filled CWC (Contingency Water Container) from the Lab humidifier. Weekly SOZh reports (on Sundays) to TsUP/Moscow deal with number & dates of water and urine containers, counter readings of water consumption & urine collection, plus data and total operating time of the Russian POTOK-150MK (150 micron) air filter unit of the SOGS air revitalization subsystem.)

Eyharts collected a fluid sample from the Columbus module ITCS (Internal Thermal Control System) and transferred it to the Shuttle for return for ground testing & analysis.

Meanwhile, Garrett worked on the MSG (Microgravity Glovebox) rack which Peggy & Leo had relocated yesterday from the US Lab to Columbus (loc. COL1F2), today mating its power, data & fluid connectors and activating the rack.

Reisman also updated the SODF (Station Operations Data File) Emergency Books with new crew procedural instructions delivered on STS-123.

In support of last night's EVA-5, after Johnson & Reisman powered up their robotarms, the SSRMS (Space Station Remote Manipulator System) grappled the OBSS from the SRMS (Shuttle RMS) and maneuvered it to the OBSS stow position, assisting the EVA crew in attaching the OBSS to the S0/S1 truss. (SSRMS performance was nominal throughout the stow activities. This was the last planned MSS (Mobile Servicing System) operations for the docked mission.)

After return and ingress of Bob & Mike from EVA-5 last night at 10:36pm, CDR Whitson took charge of the usual post-EVA activities, i.e., photographing the EMU/spacesuit gloves and overgloves while still pressurized, recharging the EMUs with water from PWR (Payload Water Reservoir), then reconnecting the LTAs (Lower Torso Assemblies) to the EMUs and capping the UIA (Umbilical Interface Assembly), initiating the discharge/recharge process on the EMU batteries in the BSA (Battery Stowage Assembly), turning around the DCS-760 EVA camera (including downloading its EVSA and glove photographs), and reconfiguring EVA tools. (Regeneration of the METOX (Metal Oxide) CO2 filter canisters was deferred to tonight after the duct installation because of the ongoing CO2 troubleshooting.)

The crewmembers completed most of their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (FE-1, FE-2, FE-2-16), RED resistive exercise device (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Afterwards, Garrett transferred the crew's exercise data file to the MEC (Medical Equipment Computer) laptop for downlinking, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week). (For Eyharts, it was the last exercise session onboard the station.)

Sleeptime for the ISS crew began this morning at 4:00am, for the Shuttle crew at 4:30am. For both crews workday started today at 12:30pm EDT.

Main activities of today's workday (FD14): Off-duty time, PK-3 (Malenchenko), handover activities (Eyharts/Reisman), joint crews news conference (11:18pm-11:58pm); joint crews photo shoot.

Transfer Status: No concern with middeck transfers - they are back on schedule. Overall transfer is 90% complete.

STS-123 Update: TPS inspection analysis is complete and Shuttle is Go for reentry on Wednesday.

ISS Crew Sleep Shift Planning: To synchronize the ISS crew's timeline with STS-123/1J/A docking and subsequent docked activities, Peggy's, Yuri's and Leo's wake/sleep cycle underwent a number of shifts which started on 3/11. For the next three days, the wake/sleep shift schedule is as follows, getting the ISS crew back on their regular schedule by Wednesday (all times EDT):

FD14 Wake: 12:30pm (3/23) - 3:15am (3/24)
Sleep: 3:15am - 11:45am (3/24)

FD15
Wake: 11:45am (3/24) - 11:00pm (3/24)
Sleep: 11:00pm - 7:30am (3/25)

FD16
Wake: 7:30am (3/25) - 5:30pm (3/25)
Sleep: 5:30pm - 2:00am (3/26)

No CEO photo targets uplinked for today.

More... - Chronology...

Home - Browse - Contact
© / Conditions for Use